write a fraction expression equivalent to 2/3 with a denominator of 12. I know the answer, i dont how to come about the answer. please help

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

write a fraction expression equivalent to 2/3 with a denominator of 12. I know the answer, i dont how to come about the answer.
--------------------------
Ask yourself, "What would I need to multiply 3 by to get 12 in
the denominator?".
---
The answer would be "4".
----
So you need to multiply numerator and denominator by 4 to get:
---
(2/3)(4/4) = 8/12
==========================
Cheers,
Stan H.

Answer by richwmiller(17219)

so what answer did you find in the book?
2/3=x/12
solve for x
cross multiply OR
multiply both sides by 12 and both sides by 3
24=3x
8=x